the average of a class of 22 students is 21 years. the average increases by 1 when the teachers age is also included. what is the age of the teacher?

Answer by Alpha Maximus(4)

The definition of an average is the ratio of the sum of a number of values to the number of values. Since we don't have the sum, let it be 'x'
But we do have the number of values which is 22. Thus,
X/22=21, x=22*21= 462. The previously unknown sum is 462. We're now told that the average undergoes a unit increment (21+1=22), when the teachers age is added to the sum. Let the teacher's age be 'y'. Note that the number of values also increases by one(I.e: 22+1=23). Thus we have:
462+y/22=23, 462+y=506, y=506-462,y=44
Thus,it has been established that the teacher is 44yrs old.